OpenStudy (anonymous):

What is the solution to the equation? c + (–2) = –0.7 c =

OpenStudy (asad786):

1.03

OpenStudy (asad786):

wait thats not correct

OpenStudy (asad786):

1.93

OpenStudy (anonymous):

c+ (-2)=-.07 add +2 for both sides c-2+2=-.07+2 c=1.93